OpenStudy (anonymous):

1. Solve for x: −3|2x + 6| = −12 (1 point)

OpenStudy (anonymous):

First isolate the inequality. Can you do that?

OpenStudy (anonymous):

How do you get the \(-3\) to the other side?

OpenStudy (anonymous):

dropping the x

OpenStudy (anonymous):

No, you want to divide by \(-3\)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[ |2x+6| =4 \]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now you have two equations: \[ 2x+6=4 \]And \[ -(2x+6)=4 \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You need to solve for \(x\) in each equation. Can you do it?

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

Good, there are two answers.

OpenStudy (anonymous):

can you tell me what they are

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%E2%88%923%7C2x+%2B+6%7C+%3D+%E2%88%9212 Under "Integer Solutions" But you should do it yourself.

OpenStudy (anonymous):

okay thanks